write as an equation and then solve
15 is 3 less than half a number
Im thinking of a number. Ill subtract 7 and then multiply the result by 3, I get 18. What was the original number?
write in standard form: y=3x-1
whats the slope of (-5,2) and (1,4)
write the slope-intercept form of 3x+y=4
find slope of 3x+y=4

Answer by DJ_Sammie(24)

let n = the number
15=1/2n-3

let n = the number
(n-7)*3=18
n=6
2-4/-5-1=m
m=slope
so -2/-6=m
1/3=m

3x+y=4
-3x -3x
y=-3x+4

3x+y=4
-3x -3x
y=-3x+4
y=mx+b
m=slope
so the slope is -3